OpenStudy (anonymous):

Hello, I need help to solve this ϕ(x)=λ∫2π0(cos2(x+y)+1/2)ϕ(y)dy, The eigenvectors and eigenvalues...

OpenStudy (experimentx):

try differentiating twice with respect to x, it seems that it will satisfy a DE of order 2.

OpenStudy (anonymous):

That's not the method I use

OpenStudy (experimentx):

what method are you using then?

OpenStudy (experimentx):

\[ (\cos^2(x+y) + 1/2) = \frac 1 2 \cos 2(x+y) \]